# On the symmetry energy expansion and the peak value of the bulk viscosity

### v1

@@ Abstract
The symmetry energy expansion is a useful way to parametrize the properties of dense matter near nuclear saturation density, and much work has been done to connect physical quantities like the neutron star radius and the core-crust transition density to the symmetry energy parameters. In this work, I connect the weak-interaction-driven bulk viscosity in neutron-proton-electron ($npe$) matter to the symmetry parameters by calculating the susceptibilities of dense matter in terms of the symmetry energy. I use this result to calculate the resonant-peak value of the bulk viscosity as a function of density, finding that it strongly depends on $L$, as does the minimum bulk-viscous dissipation timescale. Also resulting from this calculation is a formula for finding the conformal points of the zero-temperature equation of state. Finally, I determine for which values of the symmetry parameters the maximum r-mode stable rotation frequency of an $npe$-matter neutron star is smaller than the Kepler frequency, in the high-temperature conditions where bulk viscosity is the dominant dissipation mechanism.
@@ I Introduction
The equation of state (EoS) of dense matter describes thermodynamic properties of the matter, such as the energy density $\varepsilon$ and pressure $P$, as a function of its baryon density $n_{B}$ and temperature $T$ [1]. The EoS inherently applies only to matter that is in thermal equilibrium, but the assumption of thermal equilibrium is a good one for matter in neutron stars, neutron star mergers, and supernovae (with the exception of, in some conditions, neutrinos [2, 3, 4] and perhaps feebly interacting particles representing physics beyond the standard model [5, 6, 7]). Neutron star matter is multicomponent, containing (at the very least) neutrons, protons, and electrons, and while weak interactions attempt to drive the matter towards beta (chemical) equilibrium, dense matter is likely to be out of beta equilibrium in a wide number of situations of current interest [3, 8, 9, 10].
Great effort has been undertaken to calculate and measure the EoS of dense matter, as it directly controls the allowed masses and radii of neutron stars [11, 1, 12, 13, 14], the extent to which they are tidally deformed in the inspiral phase of a neutron star merger [11, 13, 15], and the oscillation spectrum of both isolated neutron stars [16, 17] and neutron star merger remnants [18, 19, 20]. A natural starting point for calculating the EoS is nuclear saturation density, $n_{0}\equiv 0.16\text{ fm}^{-3}$, as this density is reached in atomic nuclei, providing a wealth of experimental input [11, 21]. In this approach, one Taylor expands the EoS around the conditions encountered in nuclei [nuclear saturation density and symmetric nuclear matter ($x_{p}=1/2$)] and introduces the symmetry energy: the energy difference between symmetric and pure neutron matter. The coefficients of the Taylor expansion have become the object of much research and the lower order terms are well constrained [21, 22].
@@
However, in neutron star contexts, there is often little linkage made between the EoS and transport: typically the two are calculated in entirely different frameworks. This often seems natural, because the EoS is dominated by strong interaction physics, while transport may be connected to a weak or electromagnetic interaction. Weak interactions, with their small cross sections, contribute negligibly to the EoS and electromagnetism contributes only at the lowest densities encountered in neutron stars. Some relevant rates for transport do involve strong interactions, but often the strong interaction is modeled with one-pion exchange [50], and in the mean field approximation, a common technique used in EoS construction, the pion contribution to the EoS vanishes [51]. Connecting the EoS and approaches used in transport calculations in neutron stars is not necessarily easy.
@@
In this work, I connect the resonant-peak value of the bulk viscosity of neutron-proton-electron ($npe$) matter to the symmetry parameters and explore the impact of current constraints on those coefficients on transport in dense matter. In Sec. II, I discuss the symmetry energy expansion and constraints on its coefficients. The susceptibilities of $npe$ matter are derived in Sec. III, and in Sec. IV, I derive the resonant-peak value of the bulk viscosity and use that result to calculate the minimum bulk-viscous dissipation timescale for small-amplitude density oscillations. Finally, I apply the peak bulk viscosity result to the r-mode instability window to calculate the maximum stable neutron star rotation frequency.
I work in natural units, where $\hbar=c=k_{B}=1$.
@@ II The symmetry energy expansion
Near nuclear saturation density, matter is expected to be $npe$ matter: a strongly interacting fluid of neutrons and protons with a free Fermi gas of electrons. In this article, I will focus only on the zero temperature EoS. At a given baryon density $n_{B}\equiv n_{n}+n_{p}$ the proton fraction assumes a specific value in beta equilibrium, $x_{p}^{\text{eq.}}$. However, in many astrophysical situations of relevance, the system can be out of beta equilibrium by an amount
@@
In $npe$ matter, the first law of thermodynamics is [49]
@@
The leptons are treated as a free Fermi gas and their mass is taken to be zero. The treatment of the baryonic sector is explained below.
The idea of the symmetry energy expansion is to Taylor expand the energy per baryon $(E_{\text{nuc}}/N_{B})(n_{B},x_{p})$ around a point that we understand very well, namely the conditions encountered in nuclei: $\left\{n_{B},x_{p}\right\}=\left\{n_{0},1/2\right\}$. For historical reasons, the expansion is in powers of $(n_{B}-n_{0})/(3n_{0})$ and $(1-2x_{p})$. Neutron star matter is not close to symmetric, so carrying forward with expansion around $x_{p}=1/2$ is not appropriate. It is conventional to terminate the expansion in $(1-2x_{p})$ at quadratic order, called the parabolic approximation, and to assume that this $x_{p}$-dependence encapsulates the entirety of the $x_{p}$-dependence of the baryonic part of the EoS. There is good evidence to support this approximation [64, 65, 66]. Thus, the energy per baryon in the nuclear sector is
@@
The coefficient $S(n_{B})$ of the quadratic term is called the symmetry energy. Now, both the symmetry energy and the symmetric matter expressions are Taylor expanded in density around $n_{0}$. The symmetry energy $S(n_{B})$ is given by
@@ II.1 Constraints on the Symmetry Parameters
The quantity $m-B_{\text{sat}}$ is the nucleon mass minus the binding energy in symmetric nuclear matter at $n_{0}$. With reference to the semi-empirical mass formula for nuclei, the binding energy is determined to be $B_{\text{sat}}=16\pm 1$ MeV [21, 11] (though see the critique of Atkinson et al. [69]). The incompressibility $K$ is measured from giant monopole resonances in heavy nuclei, yielding $K\approx 230\pm 20$ MeV [21]. The skewness $Q$ is essentially unconstrained: Burgio et al. [67] mention an estimate 500 $-500\text{ MeV}<Q<300\text{ MeV}$, while Li et al. [22] quote 800 $-800\text{ MeV}<Q<400\text{ MeV}$.
The symmetry energy coefficients are simultaneously constrained in the laboratory by nuclear mass data, neutron skin measurements, and measurements of collective flow in relativistic heavy ion collisions [21]. Relatively strong constraints are possible on the lower order coefficients, but the higher order coefficients are subject to vast uncertainties. The symmetry energy at saturation density, $J$, is well constrained, lying in the range 29 $29\text{ MeV}<J<34\text{ MeV}$ [11] The slope parameter $L$ is much less certain. A collection of nuclear experiments have constrained 43 $43\text{ MeV}<L<75\text{ MeV}$ [11]. Drischler et al. calculated $L$ with chiral effective field theory ($\chi$EFT) and found $L=59.8\pm 4.1$ MeV. The results of the PREX-II experiment [70] lead to the prediction of a particularly large value $L=106\pm 37$ MeV [71]. However, Essick et al. combined the PREX-II experiment results with $\chi$EFT and astrophysical data and inferred 38 $38\text{ MeV}<L<67\text{ MeV}$ (90% credible interval), which is in mild tension with the large value predicted by PREX-II [72]. A more recent calculation combining neutron star data and $\chi$EFT shows 44 $44\text{ MeV}<L<70\text{ MeV}$ (95% credible interval) [73].
As mentioned, $K_{\text{sym}}$ is even less certain. Essick et al. find 221 $-221\text{ MeV}<K_{\text{sym}}<27\text{ MeV}$ [72], while Li et al. mention a possible range 400 $-400\text{ MeV}<K_{\text{sym}}<100\text{ MeV}$ [22]. Most analyses lead to a negative $K_{\text{sym}}$, though recently a few models with large $K_{\text{sym}}$ were produced with the goal of reconciling the PREX and CREX measurements [74]. Finally, $Q_{\text{sym}}$ is virtually unknown: Li et al. suggest 200 $-200\text{ MeV}<Q_{\text{sym}}<800\text{ MeV}$ [22].
In producing an EoS, for example, one built on a Skyrme model [75] or an RMF theory [76], one introduces parameters that are then fit to experiment including properties of finite nuclei and of bulk nuclear matter. Each of these EoSs has its own predicted set of symmetry coefficients, and correlations between the symmetry coefficients are observed [21]. Fitting nuclear masses leads to a strong linear correlation between $J$ and $L$ [21]. Several correlations are listed by Li & Magno [77], including ones derived from Skyrme and RMF theories [78, 79] and from properties of Fermi liquid theory [80]. In this paper, I will use the relations given by Tews et al., [78]
@@
The linear relationships are based around a median relationship with a plus/minus offset added to encompass 95.4% of the studied EoSs (the EoSs were first screened by Tews et al. for compatibility with experiment).
@@ II.2 The EoS in terms of Symmetry Parameters
From Eq. 4, the energy density of dense matter is
@@
By setting $\delta\mu=0$, we obtain an equation for the proton fraction in beta equilibrium $x_{p}^{\text{eq.}}$
@@
The proton fraction in beta equilibrium is a function of density (c.f. Fig. 2 in [49]). An interesting potential feature in this curve is a conformal point: a point where $x_{p}^{\text{eq.}}(n_{B})$ reaches a local maximum (or minimum). Conformal points have important implications for the bulk viscosity [49, 43]. We can use Eq. 12 to find conformal points in the EoS. Differentiating Eq. 12 with respect to density, solving for $\mathop{dx_{p}^{\text{eq.}}}/\mathop{dn_{B}}$ and then setting it to zero yields the condition
@@
[…]
@@
±$\pm\sqrt{\left(2L+K_{\text{sym}}\right)^{2}-10K_{\text{sym}}\left(L-J\right)}\bigg{]}$
[…]
@@ III Bulk viscosity
[…]
@@
At fixed frequency $\omega$, the subthermal bulk viscosity is a function of density and temperature. The susceptibility prefactor, and the susceptibilities themselves, depend on density but very little on temperature (c.f. Fig. 3 in [35]). The beta-equilibration rate $\gamma$ depends strongly on temperature, due to the rapid opening up of phase space as the Fermi-Dirac distributions broaden with temperature, but relatively little on density [^2] At low temperature, the direct Urca threshold in $npe$ matter introduces a sharp change in $\gamma$ at the density where $x_{p}=1/9$. However, as temperature rises above about 1 MeV, the threshold blurs across density and even below the direct Urca threshold, direct Urca begins to dominate over modified Urca. This effect lessens the density dependence [89, 90]. (c.f. Fig. 8 in [49] or Fig. 7 in [91]).
The $\gamma/\left(\omega^{2}+\gamma^{2}\right)$ factor in the bulk viscosity has a resonance structure and peaks as $\gamma$ approaches $\omega$. Physically, this corresponds to actual value of the proton fraction $x_{p}$ being maximally out of phase with its beta-equilibrium value $x_{p}^{\text{eq.}}$, which leads to maximal $PdV$ work done and thus maximal bulk-viscous damping [49]. To visualize the behavior of the bulk viscosity in the density-temperature plane, one relies on the fact that the susceptibility prefactor is mostly a function of density and the resonance factor is mostly a function of temperature. At fixed density, the bulk viscosity is sharply peaked as a function of temperature, with the peak occuring at the temperature where $\gamma=\omega$. The height of the peak is set, in part, by the susceptibility prefactor. At the next density, the bulk viscosity as a function of temperature is peaked at roughly the same temperature, but the peak height may be different if $A^{2}/B$ has changed significantly. Thus, the bulk viscosity has a ridge-like structure in density-temperature plan. Contour plots of the bulk viscosity in the density-temperature plane are displayed in Fig. 4 in both [35] and [44]. Also of interest is Fig. 2 in [37], where the bulk viscosity is mapped on to a neutron star merger remnant.
@@
$\left|\dfrac{A^{2}}{B}\right|=\dfrac{4n_{0}}{3}x_{p}^{\text{eq.}}\left[1-8x_{p}^{\text{eq.}}+36(x_{p}^{\text{eq.}})^{2}\right]\bigg{\{}\dfrac{\left(L-J\right)^{2}}{J}-\dfrac{L-J}{3J^{2}}\left(3J^{2}-2JK_{\text{sym}}-8JL+L^{2}\right)\left(\dfrac{n_{B}-n_{0}}{n_{B}}\right)$ (20)

### v3

@@ Abstract
The symmetry energy expansion is a useful way to parametrize the properties of dense matter near nuclear saturation density, and much work has been done to connect physical quantities like the neutron star radius and the core-crust transition density to the symmetry energy parameters. In this work, I connect the weak-interaction-driven bulk viscosity in neutron-proton-electron (npenpe) matter to the symmetry parameters by calculating the susceptibilities of dense matter in terms of the symmetry energy. I use this result to calculate the resonant-peak value of the bulk viscosity as a function of density, finding that it strongly depends on $L$, as does the minimum bulk-viscous dissipation timescale. Also resulting from this calculation is a formula for finding the conformal points of the zero-temperature equation of state. Finally, I determine for which values of the symmetry parameters the maximum r-mode-stable rotation frequency of an npenpe-matter neutron star is smaller than the Kepler frequency, in the high-temperature conditions where bulk viscosity is the dominant dissipation mechanism.
@@ I Introduction
The equation of state (EoS) of dense matter describes thermodynamic properties of the matter, such as the energy density $\varepsilon$ and pressure $P$, as a function of its baryon density $n_{B}$ and temperature $T$ [1]. The EoS inherently applies only to matter that is in thermal equilibrium, but the assumption of thermal equilibrium is a good one for matter in neutron stars, neutron star mergers, and supernovae (with the exception of, in some conditions, neutrinos [2, 3, 4] and perhaps feebly-interacting particles representing physics beyond the standard model [5, 6, 7]). Neutron star matter is multicomponent, containing (at the very least) neutrons, protons, and electrons, and while weak interactions attempt to drive the matter towards beta (chemical) equilibrium, dense matter is likely to be out of beta equilibrium in a wide number of situations of current interest [3, 8, 9, 10].
Great effort has been undertaken to calculate and measure the EoS of dense matter, as it directly controls the allowed masses and radii of neutron stars [11, 1, 12, 13, 14], the extent to which they are tidally deformed in the inspiral phase of a neutron star merger [11, 13, 15], and the oscillation spectrum of both isolated neutron stars [16, 17] and neutron star merger remnants [18, 19, 20]. A natural starting point for calculating the EoS is nuclear saturation density, $n_{0}\equiv 0.16\text{ fm}^{-3}$, as this density is reached in atomic nuclei, providing a wealth of experimental input [11, 21]. In this approach, one Taylor expands the EoS around the conditions encountered in nuclei [nuclear saturation density and symmetric nuclear matter ($x_{p}=1/2$)] and introduces the symmetry energy: the energy difference between symmetric and pure neutron matter. The coefficients of the Taylor expansion have become the object of much research and the lower order terms are fairly well constrained [21, 22].
@@
However, in neutron star contexts there is often little linkage made between the EoS and transport: typically the two are calculated in entirely different frameworks. This often seems natural, because the EoS is dominated by strong interaction physics, while transport may be connected to a weak or electromagnetic interaction. Weak interactions, with their small cross sections, contribute negligibly to the EoS and electromagnetism contributes only at the lowest densities encountered in neutron stars. Some relevant rates for transport do involve strong interactions, but often the strong interaction is modeled with one-pion exchange [50] and in the mean field approximation (a common technique used in EoS construction) the pion contribution to the EoS vanishes [51]. Connecting the EoS and approaches used in transport calculations in neutron stars is not necessarily easy.
@@
In this work, I connect the resonant-peak value of the bulk viscosity of neutron-proton-electron (npenpe) matter to the symmetry parameters and explore the impact of current constraints on those coefficients on transport in dense matter. In Sec. II, I discuss the symmetry energy expansion and constraints on its coefficients. The susceptibilities of npenpe matter are derived in Sec. III, and in Sec. IV, I derive the resonant-peak value of the bulk viscosity and use that result to calculate the minimum bulk-viscous dissipation timescale for small-amplitude density oscillations. Finally, I apply the peak bulk viscosity result to the r-mode instability window to calculate the maximum neutron star rotation frequency that is stable to r-modes.
I work in natural units, where $\hbar=c=k_{B}=1$. All data presented in the figures can be found in the Zenodo repository [65].
@@ II The symmetry energy expansion
Near nuclear saturation density, matter is expected to be npenpe matter: a strongly interacting fluid of neutrons and protons with a free Fermi gas of electrons. In this article, I will focus only on the zero temperature EoS. At a given baryon density $n_{B}\equiv n_{n}+n_{p}$ the proton fraction $x_{p}\equiv n_{p}/n_{B}$ assumes a specific value $x_{p}^{\text{eq.}}$ in beta equilibrium. However, in many astrophysical situations of relevance, the system can be out of beta equilibrium by an amount
@@
In zero-temperature npenpe matter, the first law of thermodynamics is [49]
@@
The leptons (electrons) are treated as a free Fermi gas and their mass is taken to be zero. The treatment of the baryonic sector is explained below.
The idea of the symmetry energy expansion is to Taylor expand the energy per baryon $E_{\text{nuc}}/N_{B}$ around a point that is understood very well, namely the conditions encountered in nuclei: $\left\{n_{B},x_{p}\right\}=\left\{n_{0},1/2\right\}$. For historical reasons, the expansion is in powers of $(n_{B}-n_{0})/(3n_{0})$ and $(1-2x_{p})$. Neutron star matter is not close to symmetric, so carrying forward with the expansion around $x_{p}=1/2$ is not appropriate. It is conventional to terminate the expansion in $(1-2x_{p})$ at quadratic order and to assume that this $x_{p}$-dependence encapsulates the entirety of the $x_{p}$-dependence of the baryonic part of the EoS. This is called the parabolic approximation and there is good evidence to support its approximate validity [64, 66, 67]. Thus, the energy per baryon in the nuclear sector is
@@
The coefficient $S(n_{B})$ of the quadratic term is called the symmetry energy. Now, both the symmetry energy and the symmetric matter expressions are Taylor expanded in density around $n_{0}$. The symmetry energy is given by
@@ II.1 Constraints on the Symmetry Parameters
The quantity $m-B_{\text{sat}}$ is the nucleon mass (in vacuum) minus the binding energy in symmetric nuclear matter at $n_{0}$. With reference to the semi-empirical mass formula for nuclei, the binding energy is determined to be $B_{\text{sat}}=16\pm 1$ MeV [21, 11] (though see the critique of Atkinson et al. [70]). The incompressibility $K$ is measured from giant monopole resonances in heavy nuclei, yielding $K\approx 230\pm 20$ MeV [21]. The skewness $Q$ is essentially unconstrained: Burgio et al. [68] mention an estimate $500\text{ MeV}<Q<300\text{ MeV}$, while Li et al. [22] quote $800\text{ MeV}<Q<400\text{ MeV}$.
[…]
As mentioned, $K_{\text{sym}}$ is even less certain. Essick et al. find $221\text{ MeV}<K_{\text{sym}}<27\text{ MeV}$ [74], while Li et al. mention a possible range $400\text{ MeV}<K_{\text{sym}}<100\text{ MeV}$ [22]. Most analyses lead to a negative $K_{\text{sym}}$, though recently a few models with large $K_{\text{sym}}$ were produced with the goal of reconciling the PREX and CREX measurements [78]. Finally, $Q_{\text{sym}}$ is virtually unknown: Li et al. suggest 200 $0\text{ MeV}<Q_{\text{sym}}<800\text{ MeV} [2$2].
In producing an EoS, for example, one built on a Skyrme model [79] or an RMF theory [80], one introduces parameters that are then fit to experiment including properties of finite nuclei and of bulk nuclear matter [51]. Each of these EoSs has its own predicted set of symmetry coefficients, and correlations between the symmetry coefficients are observed [21]. Fitting nuclear masses leads to a strong linear correlation between $J$ and $L$ [21]. Several correlations are listed by Li & Magno [81], including ones derived from Skyrme and RMF theories [82, 83] and from properties of Fermi liquid theory [84]. In this paper, I will use the relations given by Tews et al. [82]
@@
The linear relationships are based around a median relationship with a plus/minus offset added to encompass 95.4% of the studied EoSs (the EoSs were first screened by Tews et al. for compatibility with experiment). Similar linear relationships were seen in a different class of EoSs developed by Oyamatsu [85].
@@ II.2 The EoS in terms of Symmetry Parameters
From Eq. 4, the energy density of dense matter, not necessarily in beta equilibrium, is
@@
By setting [^1] At finite temperature, if the matter remains neutrino-transparent, this beta-equilibrium condition picks up finite-temperature corrections [86, 87, 44, 8], but I ignore these here. $\delta\mu=0$, one obtains an equation for the proton fraction in beta equilibrium $x_{p}^{\text{eq.}}$
@@
The proton fraction in beta equilibrium is a function of density (c.f. Fig. 2 in [49]). An interesting potential feature in this curve is a conformal point: a point where $x_{p}^{\text{eq.}}(n_{B})$ reaches a local maximum (or minimum). Conformal points have important implications for the bulk viscosity [49, 43]. Eq. 12 can be used to find conformal points in the EoS. Differentiating Eq. 12 with respect to density, solving for $\mathop{dx_{p}^{\text{eq.}}}/\mathop{dn_{B}}$ and then setting it to zero yields the condition
@@
$\alpha^{\text{conf.}}=-\dfrac{3}{2}\left(\dfrac{L-J}{L}\right)\quad\text{[linear order in $S(n_{B})$]}$ (14)
@@
$\pm\sqrt{\left(2L+K_{\text{sym}}\right)^{2}-10K_{\text{sym}}\left(L-J\right)}\bigg]$
$\text{[quadratic order in $S(n_{B})$]}.$
@@ III Bulk viscosity
The combination of susceptibilities in the prefactor of the bulk viscosity turns out to be related to the difference between the adiabatic and equilibrium incompressibilities of dense matter [^2] BB is negative, so the adiabatic (fixed $x_{p}$) incompressibility is larger than the equilibrium (fixed $\delta\mu$) incompressibility. [43, 49, 93, 94, 95]
@@
At fixed frequency $\omega$, the subthermal bulk viscosity is a function of density and temperature. The susceptibility prefactor, and the susceptibilities themselves, depend on density but very little on temperature (c.f. Fig. 3 in [35]). The beta-equilibration rate $\gamma$ depends strongly on temperature, due to the rapid opening up of phase space as the Fermi-Dirac distributions broaden with temperature, but depends relatively little on density [^3] At low temperature, the direct Urca threshold in npenpe matter introduces a sharp change in $\gamma$ at the density where $x_{p}=1/9$. However, as temperature rises above about 1 MeV, the threshold blurs across density and even below the direct Urca threshold, direct Urca begins to dominate over modified Urca. This effect lessens the density dependence [86, 87]. (c.f. Fig. 8 in [49] or Fig. 7 in [96]).
The $\gamma/\left(\omega^{2}+\gamma^{2}\right)$ factor in the bulk viscosity has a resonance structure and peaks as $\gamma$ approaches $\omega$. Physically, this corresponds to actual value of the proton fraction $x_{p}$ being maximally out of phase with its beta-equilibrium value $x_{p}^{\text{eq.}}$, which leads to maximal $PdV$ work done and thus maximal bulk-viscous damping [49]. To visualize the behavior of the bulk viscosity in the density-temperature plane, one relies on the fact that the susceptibility prefactor is mostly a function of density and the resonance factor is mostly a function of temperature. At fixed density, the bulk viscosity is sharply peaked as a function of temperature, with the peak occurring at the temperature where $\gamma=\omega$. The height of the peak is set, in part, by the susceptibility prefactor. At the next density, the bulk viscosity as a function of temperature is peaked at roughly the same temperature, but the peak height may be different if $A^{2}/B$ has changed significantly. Thus, the bulk viscosity has a ridge-like structure in density-temperature plan. Contour plots of the bulk viscosity in the density-temperature plane are displayed in Fig. 4 in both [35] and [44]. Also of interest is Fig. 2 in [37], where the bulk viscosity is mapped on to a neutron star merger remnant.
@@
$\left|\dfrac{A^{2}}{B}\right|=\dfrac{4n_{0}}{3}x_{p}^{\text{eq.}}\left[1-8x_{p}^{\text{eq.}}+36(x_{p}^{\text{eq.}})^{2}\right]\bigg\{\dfrac{\left(L-J\right)^{2}}{J}-\dfrac{L-J}{3J^{2}}\left(3J^{2}-2JK_{\text{sym}}-8JL+L^{2}\right)\left(\dfrac{n_{B}-n_{0}}{n_{0}}\right)$ (20)